OpenStudy (anonymous):

PLEASE HELP I WILL DAN AND MEDAL!!! Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

HanAkoSolo (jamierox4ev3r):

let x represent girls and y represent boys

OpenStudy (anonymous):

89=g2+8+b....?

OpenStudy (campbell_st):

well the sum or boys and girls in sport is b + g = 89 and for numbers boys to girls b = 2g + 8 which can be written as b - 2g = 8 so you have your 2 equations b + g = 78 b - 2g = 8 an easy solution is double the 1st equation and add to the 2nd equation, which will eliminate g. then you can solve for b. when you know the number of boys, substitute that number into the 1st equation to find the number of girls. hope this makes sense

OpenStudy (anonymous):

b+g=89, b=2g+8, g=27 b=62

OpenStudy (campbell_st):

lol... why worry about learning when people just give answers.